# Hydrostatic mass of galaxy clusters within some theories of gravity

### v2

@@ Abstract
The mass of galaxy clusters (GCs) can be determined by calculating the hydrostatic equilibrium equation. In this work, we revisit the derivation of hydrostatic mass of GCs within modified emergent Newtonian gravity (MENG) with generalized uncertainty principle (GUP) correction. We also derive the hydrostatic mass of GCs within Eddington-inspired Born-Infeld (EiBI) theory, and beyond Horndeski gravity (BHG). We apply the formulations on the masses of 10 GCs. We compare our results with the Newtonian mass of GCs. For a regime, we get an insight that all formulations could match the Newtonian mass. Thus, the impact of the modified theories of gravity used in this work can be neglected in this regime. The noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG. We also compare our results with the GCs’ baryonic masses $M_{bar}$. In this comparison, only the MENG formulation with $\beta_{0}=2.70\times 10^{115}$ produces the best linear fit with the slope $\frac{M_{MENG}}{M_{bar}}=0.755\pm 0.061$. This means that in this work, MENG is the best theory of gravity that alleviates the mass discrepancy of hydrostatic mass and baryonic mass of GCs.
@@
hydrostatic mass , galaxy cluster , emergent gravity , Eddington-Inspired Born-Infeld theory , beyond Horndeski gravity
@@ 1 Introduction
Another attempt to obtain the mass of GCs was achieved by Apryandi et. al. [17]. Using the generalized uncertainty principle (GUP) and following Verlinde’s emergent gravity or entropic gravity procedure [18] to get Newtonian gravity formula, they derived the modified emergent Newtonian gravity (MENG). The formalism was then used to derive a formula to calculate the mass of GCs. Unfortunately, in their work, they have not yet applied their formula to calculate GCs’ masses, like what has been done by Desai [16] and Gupta & Desai [3].
Basically, Verlinde’s theory of emergent gravity describes gravity as an emergent phenomenon rather than a fundamental force. One can impose this reasoning on de Sitter space to find that gravity behaves differently on galaxy and GC scales. This excess gravity might offer an alternative to dark matter (DM) [19]. The emergent gravity paradigm is such an attempt to obtain and interpret the field equations of gravitational theories in a thermodynamic context [20]. Please see Refs. [18, 19, 20] for further discussions on emergent gravity.
Unfortunately, Tamosiunas et al. [19] found that GR is preferred over emergent gravity in all tested datasets of galaxies and GCs. They discovered that the simultaneous emergent gravity fits of the X-ray and weak lensing datasets are significantly worse than those provided by GR with cold dark matter. For the Coma cluster, the predictions from emergent gravity and GR agree in the range of 250--700 kpc. In contrast, at around 1 Mpc scale, emergent gravity total mass predictions are larger by a factor of 2. For the cluster stack the predictions are only in good agreement at around the 1--2 Mpc scales, while for the radius $r\gtrsim$ 10 Mpc, emergent gravity is in strong tension with the data. It is important to keep in mind that GUP’s effect is not included in the work of Tamosiunas et al. Although controversial [21, 22, 23], Verlinde’s theory of emergent gravity has good agreement with some observations (please see Refs. [24, 25, 26, 27]).
@@
An example of the modified theory of gravity is the Eddington-inspired Born-Infeld (EiBI) theory. Santos & Santos [35] introduced the geometrical mass of GCs within EiBI theory. The main problem they pointed out is the mass discrepancy in GCs, which is usually explained by a postulate of the existence of a cold pressureless DM. Unfortunately, the nature of DM is still unknown and the only existing evidence for its presence is gravitational. Thus, another possible way to overcome the problem of discrepancy of the GC’s mass is by modifying the theory of gravity. In their work, they derived the virial theorem in EiBI theory and showed that the geometrical mass, which is a mathematical artifact in the theorem, may account for the virial mass discrepancy in GCs.
@@
Motivated by the previous studies, in this work, we calculate the hydrostatic mass of several GCs presented in Ref. [3] using MENG formulation derived by Apryandi et. al. [17]. Besides that, we also derive the formulation for the hydrostatic mass of GCs within EiBI theory and BHG and then obtain the EiBI and BHG version of numerical calculation on the hydrostatic mass of the GCs which is previously numerically calculated using MENG. Finally, the results are compared with the corresponding Newtonian values and with the baryonic mass $M_{bar}$ of the GCs presented in Ref. [41]. We expect that the baryonic mass and the hydrostatic mass would essentially agree if the free parameters of MENG, EiBI, and BHG were correctly represented by our formulation.
This paper is organized as follows. In section 2, we explain the construction of formalism we use, where the MENG formulation derived in Ref. [17] is revisited, and the derivations of hydrostatic mass of GCs within EiBI theory and BHG are presented. In section 3, we present and discuss the numerical results of GC masses calculation. Lastly, section 4 contains the concluding remarks of this work.
@@ 2 Formalisms
To make the discussion self-contained, in subsection 2.1, we revisit the hydrostatic mass of GCs within MENG formalism derived by Apryandi et al. [17]; in subsection 2.2 we derive the hydrostatic mass of GCs within EiBI theory; while derivation for the hydrostatic mass of GCs within BHG is given in subsection 2.3.
2.1 Hydrostatic mass of GCs within MENG
In this subsection, we start with the GUP. It is worth keeping in mind the ordinary unmodified Heisenberg uncertainty, i.e. $\Delta x\Delta p\geq\frac{\hbar}{2}$. Here we use the quadratic GUP, which reads
$\Delta x\Delta p\geq\frac{\hbar}{2}\left(1+\beta p^{2}\right),$ (1)
where $\beta=\frac{\beta_{0}l_{p}^{2}}{\hbar^{2}}$, with $\beta_{0}$ is a free dimensionless parameter of GUP, and Planck length $l_{p}=\sqrt{\frac{G\hbar}{c^{3}}}\approx$ 10-35 m. Moreover, the modified minimum volume of phase-space writes
$\Delta V_{GUP}=\hbar^{3}\left(1+\beta p^{2}\right)^{3}.$ (2)
Due to Eq. (2), the volume invariance of phase-space has to be transformed, i.e.
$\int\frac{d^{3}x\>d^{3}p}{\hbar^{3}}\longrightarrow\int\frac{d^{3}x\>d^{3}p}{\hbar^{3}\left(1+\beta p^{2}\right)^{3}}.$ (3)
Based on the transformed volume invariance showed by Eq. (3), we can write the expression of energy surface equation, i.e.
$\sum E=\frac{1}{h^{3N}}\int\frac{d^{3}x_{1}...d^{3}x_{N}\>d^{3}p_{1}...d^{3}p_{N}}{(1+\beta p^{2})^{3N}}.$ (4)
Here $N$ denotes the number of particles that are related to the equation of state (EOS) of an ideal gas, i.e. $PV=Nk_{b}T$, where $P$ is pressure, $V$ is volume, $k_{b}$ is Boltzmann constant, and $T$ is the absolute temperature. The EOS of ideal gas becomes very important in this study since we follow the procedure in Ref. [3] in using the EOS of ideal gas as the EOS of GCs.
By applying $p^{2}=2m_{b}E$, where $m_{b}$ denotes the mass of an individual atom of the ideal gas; and the entropy $S=k_{b}\ln(\sum E)$, it is easy to show that
$S=Nk_{b}\left[\frac{5}{2}+\ln\left\{\left(\frac{V}{N\hbar^{3}(1+2\beta m_{b}E)^{3}}\right)\left(\frac{4\pi m_{b}E}{3N}\right)^{3/2}\right\}\right].$ (5)
The absolute temperature $T=\frac{\partial E}{\partial S}$ now becomes
$T=\left(\frac{1+2\beta m_{b}E}{1-2\beta m_{b}E}\right)\left(\frac{2E}{3k_{b}N}\right);$ (6)
while the pressure $P=-\frac{\partial E}{\partial V}$ writes
$P=\left(\frac{1+2\beta m_{b}E}{1-2\beta m_{b}E}\right)\frac{2E}{3V}.$ (7)
On the other hand, a system of bits on the holographic surface satisfies the ideal gas distribution. The amount of information or the amount of bits on the holographic surface is given by [18]
$N=\frac{Ac^{3}}{G\hbar},$ (8)
where $A$ denotes the surface area, and $c$ denotes the speed of light. Considering Bekenstein-Hawking entropy [42, 43] $S_{BH}=\frac{k_{b}A}{4l_{p}^{2}}$, Eq. (8) now reads
$N=\frac{4S_{BH}}{k_{b}}.$ (9)
After involving GUP, Eq. (9) is modified to be
$\mathcal{N}=\frac{A}{l_{p}^{2}}-\frac{\pi\beta_{0}}{4}\ln\left(\frac{A}{4l_{p}^{2}}\right);$ (10)
while Bekenstein-Hawking entropy is modified to be
$\mathcal{S}_{BH}=k_{b}\frac{A}{4l_{p}^{2}}-\frac{k_{b}\pi\beta_{0}}{16}\ln\left(\frac{A}{4l_{p}^{2}}\right).$ (11)
Verlinde’s entropic gravity reads [18]
$F=T\frac{\Delta S}{\Delta x}.$ (12)
By involving the modified equations to Verlinde’s procedure, we obtain the formula of modified emergent Newtonian gravitational force, i.e.
$\mathcal{F}=G\frac{Mm}{r^{2}}\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right].$ (13)
As we expect, if the GUP recovers to ordinary Heisenberg uncertainty principle (i.e. $\beta=0$), then Eq. (13) recovers to the ordinary familiar formula of Newtonian gravitational force, $F=G\frac{Mm}{r^{2}}$.
The equation of hydrostatic equilibrium now writes
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right],$ (14)
and for the ideal gas,
where $\mu$ is the mean molecular weight of the cluster, and $m_{p}$ is the mass of the proton. Finally, we can easily obtain the hydrostatic mass of GCs within MENG
$M_{MENG}(r)$ $=$ $-\frac{k_{b}\>T\>r}{G\>\mu\>m_{p}}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)$ (16)
×$\times\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right].$
@@ 2.2 Hydrostatic mass of GCs within EiBI theory
Before we proceed to the derivation of GCs’ mass within EiBI theory, it is important to clarify that we involve the GUP correction on absolute temperature, which we have obtained in the subsection 2.1, in the process of the derivation. Firstly, we look at the non-relativistic limit of EiBI theory which is shown by the following Poisson equation [44, 45]
@@
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}-\frac{\kappa}{4}\>G\rho\>\frac{d\rho}{dr}.$ (18)
Now recall the equation of hydrostatic equilibrium of ideal gas shown by Eq, (15).
@@ 3 Numerical results
Figure 1: The linear fit of the relation between hydrostatic masses and the obseved baryonic mass $M_{bar}$ of 10 GCs. (a) $M_{MENG}$ with $\beta_{0}=2.70\times 10^{115}$. (b) $M_{EiBI}$ with $\kappa=5.80\times 10^{40}$. (c) $M_{BHG}$ with $\Upsilon=1.65\times 10^{71}$.
@@
(kpc) ($\mathscr{M}_{\odot}$) $\beta_{0}=-1.656\times 10^{110}$ $\kappa=5\times 10^{40}$ $\Upsilon=-0,1655\times 10^{69}$
@@
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [30], [45], and [36]. In Ref. [30], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [45], $\kappa=0,2$ could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [30], polytropic star in Ref. [45], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass.
Our result leads to an argument that if the values of the free parameters $\beta_{0}$, $\kappa$, and $\Upsilon$ are closer to zero than the values shown in Table 1, one can safely use the Newtonian calculation for hydrostatic mass of GCs, and the corrections of the gravitational theories used in this paper are negligible.
@@
For the comparison of $M_{MENG}$ with $M_{bar}$, the linear fit is shown in Fig. 1(a). The slope of this linear fit is $0.755\pm 0.061$, which is in general agreement with the MENG given the presence of various uncertainties in the estimation of $M_{MENG}$ and $M_{bar}$ of the GCs. According to what we found in this work, the value of $\beta_{0}=2.70\times 10^{115}$ in this linear fit is the finest value to reach the finest slope that is close to unity.
On the other side, the linear fit of the EiBI theory with $\kappa=5.80\times 10^{40}$ gives the slope of $0.126\pm 0.086$, while the BHG with $\Upsilon=1.65\times 10^{71}$ gives the slope of $0.126\pm 0.086$. The linear fit of EiBI theory is shown in Fig. 1(b), while the linear fit of BHG is shown in Fig. 1(c). Those slope values are below the slope value of MENG, which means that MENG is better than EiBI theory and BHG to alleviate the mass discrepancy of the hydrostatic mass and baryonic mass of GCs, with the same assumption used in Ref. [41] that there is no dark matter. It is important to note that for both EiBI theory and BHG when we set $\kappa$ and $\Upsilon$ as positive numbers, the mass of GCs decreases. As we magnify both $\kappa$ and $\Upsilon$, the hydrostatic mass of GCs increases. Based on this mathematical fact, a trivial thought might arise that we should magnify the values of $\kappa$ and $\Upsilon$ to obtain better linear fits. But in fact, it will not happen, since if we set $\kappa$ and $\Upsilon$ exceed the values used in the linear fits shown by Fig. 1(b) and Fig. 1(c), some masses of GCs reach negative values, which is not physical. By standing on this fact, we get an insight that EiBI theory and BHG cannot alleviate the mass discrepancy of hydrostatic mass and baryonic mass of GCs better than MENG.

### v3

@@ Abstract
The mass of galaxy clusters (GCs) can be determined by calculating the hydrostatic equilibrium equation. In this work, we derive the hydrostatic mass of GCs within Eddington-inspired Born-Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and modified emergent Newtonian gravity (MENG) with generalized uncertainty principle (GUP) correction. We apply the formulations on the masses of 10 GCs. We compare our results with the Newtonian mass of GCs. Within a regime, we get an insight that all formulations could match the Newtonian mass. Thus, the impact of the modified theories of gravity used in this work can be neglected in this regime. The noteworthy impact starts if we set $\kappa=5\times 10^{40}$ for EiBI theory, $\Upsilon=-0.1655\times 10^{69}$ for BHG, and $\beta_{0}=-1.656\times 10^{110}$ for MENG. We also compare our results from EiBI theory and BHG with the baryonic masses $M_{bar}$ of the GCs. A better linear fit is achieved by EiBI theory with $\kappa_{0}=5.80\times 10^{40}$, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. This leads us to the fact that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.
@@
BHG , EiBI theory , emergent gravity , galaxy clusters , hydrostatic mass
@@ 1 Introduction
Another attempt to obtain the mass of GCs was achieved by Apryandi et. al. [17]. Using the generalized uncertainty principle (GUP) and following Verlinde’s emergent gravity or entropic gravity procedure [18] to get Newtonian gravity formula, they derived the modified emergent Newtonian gravity (MENG). The formalism was then used to derive a formula to calculate the mass of GCs. However, their work has not applied the formula to calculate the masses of GCs, as has been done by Desai [16] and Gupta & Desai [3]. Additionally, there are inconsistencies in the derivation of the hydrostatic equation in Ref. [17]. These issues can be addressed using the derivation presented in Ref. [19], which also considers the GUP.
Basically, Verlinde’s theory of emergent gravity describes gravity as an emergent phenomenon rather than a fundamental force. This reasoning can be applied to de Sitter space, suggesting that gravity behaves differently on galaxy and GC scales. Such excess gravity might provide an alternative to dark matter (DM) [20]. The emergent gravity paradigm aims to derive and interpret the field equations of gravitational theories within a thermodynamic framework [21]. For further discussions on emergent gravity, see Refs. [18, 20, 21].
Notably, Tamosiunas et al. [20] found that GR is preferred over emergent gravity across all tested datasets of galaxies and GCs. They discovered that the simultaneous emergent gravity fits of the X-ray and weak lensing datasets are significantly worse than those provided by GR with cold dark matter. For the Coma cluster, the predictions from emergent gravity and GR agree in the range of 250--700 kpc. In contrast, at around 1 Mpc scale, emergent gravity total mass predictions are larger by a factor of 2. For the cluster stack the predictions are only in good agreement at around the 1--2 Mpc scales, while for the radius $r\gtrsim$ 10 Mpc, emergent gravity is in strong tension with the data. It is important to keep in mind that GUP’s effect is not included in the work of Tamosiunas et al. Although controversial [22, 23, 24], Verlinde’s theory of emergent gravity has good agreement with some observations (please see Refs. [25, 26, 27, 28]).
@@
An example of a modified gravity theory is the Eddington-inspired Born-Infeld (EiBI) theory. Santos & Santos [35] introduced the concept of geometrical mass for GCs within the framework of EiBI theory. They highlighted the primary issue of mass discrepancy in GCs, which is typically addressed by postulating the existence of cold, pressureless DM. However, the nature of DM remains unknown, and the only evidence for its existence so far is gravitational. Thus, another possible way to overcome the problem of discrepancy of the GC’s mass is by modifying the theory of gravity. In their work, they derived the virial theorem in EiBI theory and showed that the geometrical mass, which is a mathematical artifact in the theorem, may account for the virial mass discrepancy in GCs.
@@
Motivated by the previous studies, in this work, we derive the formulation for the hydrostatic mass of GCs within EiBI theory, BHG, and MENG. We then calculate the hydrostatic mass of several GCs as presented in Ref. [3], allowing free parameters which yield negligible corrections. Additionally, we vary the free parameters of the EiBI theory and BHG beyond the values that yield negligible corrections compared to the Newtonian mass of GCs. The resulting values from these varied parameters are subsequently compared with the baryonic mass $M_{bar}$ of the GCs, as provided in Ref. [41]. It should be noted that the MENG results are not compared to $M_{bar}$, as the quantum corrections from the GUP are expected to have minimal impact. We anticipate that the baryonic mass and hydrostatic mass will essentially agree if the free parameters of the EiBI theory and BHG are correctly represented by our formulation.
This paper is organized as follows. In section 2, we explain the construction of formalism we use, where the derivations of hydrostatic mass of GCs within EiBI theory and BHG are presented. Moreover, the mass of GCs within MENG formulation derived in Ref. [17] is re-derived by using the formalism in Ref. [19]. In section 3, we present and discuss the numerical results of GC masses calculation. Lastly, section 4 contains the concluding remarks of this work.
@@ 2 Formalisms
To make the discussion self-contained, subsection 2.1 derives the hydrostatic mass of GCs within EiBI theory; subsection 2.2 presents the derivation within BHG; and subsection 2.3 provides the re-derivation within MENG.
@@ 2.1 Hydrostatic mass of GCs within EiBI theory
Firstly, we look at the non-relativistic limit of EiBI theory which is shown by the following Poisson equation [42, 43]
@@
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}-\frac{\kappa}{4}\>G\rho\>\frac{d\rho}{dr},$ (2)
where $P$ denotes pressure, $G$ is the universal gravitational constant, and $\rho$ denotes mass density.
Here we use the assumption that the Equation of State (EOS) of GCs is satisfied by the EOS of ideal gases. Such EOS was also used by Gupta & Desai [3]. For the ideal gas,
where $k_{b}$ denotes Boltzmann’s constant, $\mu$ is the mean molecular weight of the cluster, $m_{p}$ is the mass of the proton, and $T$ denotes the absolute temperature.
@@ 2.2 Hydrostatic mass of GCs within BHG
2.3 Hydrostatic mass of GCs within MENG
In this subsection, we derive the hydrostatic mass of GCs within MENG. The derivation of hydrostatic equilibrium equation with GUP correction is based on the one presented in Ref. [19]. It is worth keeping in mind the ordinary unmodified Heisenberg uncertainty, i.e. $\Delta x\Delta p\geq\frac{\hbar}{2}$. Here we use the quadratic GUP, which reads
$\Delta x\Delta p\geq\frac{\hbar}{2}\left(1+\beta(\Delta p)^{2}\right),$ (14)
where $\beta=\frac{\beta_{0}l_{p}^{2}}{\hbar^{2}}$, with $\beta_{0}$ is a free dimensionless parameter of GUP, and Planck length $l_{p}=\sqrt{\frac{G\hbar}{c^{3}}}\approx$ 10-35 m. The inequality shown by Eq. (14) would give
${\Delta x\geq\hbar\sqrt{\beta}.}$ (15)
It has to be noted that Verlinde’s proposal is based on Bekenstein’s formula [46] that relates the entropy and the surface area of black holes. The increase of the black hole’s area when absorbing a classical particle whose energy $E$ and size $\Delta x$ reads
${\Delta A\geq\frac{8\pi l_{p}^{2}}{\hbar c}E\Delta x\geq\frac{8\pi l_{p}^{2}}{\hbar}\Delta p\Delta x}.$ (16)
Solving Eq. (14) to first order in $\beta$ and then substituting the solution into Eq. (16) yields
${\Delta A\geq 4\pi l_{p}^{2}\left[1+\frac{\hbar^{2}\beta}{4(\Delta x)^{2}}\right].}$ (17)
Ref. [19] agreed with Ref. [47] that $\Delta x$ can be estimated as
${\Delta x\simeq\frac{4\pi GM}{c^{2}}.}$ (18)
With this estimation, the first-order lower bound for the area increase in terms of $\beta$ is obtained as
${\Delta A_{min}\simeq 4\pi l_{p}^{2}\lambda\left[1+\frac{\hbar^{2}\beta}{4\pi A}\right],}$ (19)
where $A=4\pi\left(\frac{2GM}{c^{2}}\right)^{2}$ denotes the area of the black hole’s event horizon, and $\lambda$ is a constant that will be fixed later. For entropy, the minimum increment is typically taken as one bit, such that $\Delta S_{min}=b=\ln 2$. Thus, from these results, we can write
${\frac{dS}{dA}\simeq\frac{\Delta S_{min}}{\Delta A_{min}}=\frac{b}{4\pi\lambda l_{p}^{2}\left[1+\frac{h^{2}\beta}{4\pi A}\right]}.}$ (20)
Therefore,
${\frac{dS}{dA}\simeq\frac{b}{4\lambda\pi l_{p}^{2}}\left[1-\frac{h^{2}\beta}{4\pi A}\right].}$ (21)
Integrating Eq. (21) would give
${S(A,\beta)=\frac{b}{4\lambda\pi l_{p}^{2}}\left[A-\frac{\hbar^{2}\beta}{4\pi}\ln\left(\frac{A}{4l_{p}^{2}}\right)\right].}$ (22)
If we take a limit $\beta\rightarrow 0$, Eq. (22) is reduced to the standard Bekenstein-Hawking entropy $S_{BH}$. It can be identified that
${\frac{b}{4\pi\lambda l_{p}^{2}}=\frac{k_{b}c^{3}}{4G\hbar}.}$ (23)
Eq. (22) now writes
${S(A,\beta)=\frac{k_{b}c^{3}}{4G\hbar}\left[A-\frac{\hbar^{2}\beta}{4\pi}\ln\left(\frac{A}{4l_{p}^{2}}\right)\right].}$ (24)
Please note that Verlinde’s idea relies on the holographic principle which states that microscopic informations (i.e. the bits) of a closed surface are embedded on the surface on radius $r$. The amount of bits is given by
${N=\frac{4S}{k_{b}}}.$ (25)
In Ref. [18], Verlinde conjectured that the bits carried the energy of the system and the energy is divided evenly over the bits. The equipartition theorem gives the temperature
${T=\frac{2E}{N\>k_{b}},}$ (26)
where $E=Mc^{2}$ denotes the total energy of the system.
Following Bekenstein’s argument, as a test particle of mass $m$ approaches the surface by an amount $\Delta x$, the corresponding change in entropy associated with the boundary information reads
${\Delta S=2\pi k_{b}\frac{\Delta x}{\hbar}mc.}$ (27)
Recall the relation of $\Delta S$ to entropic force
${F\Delta x=T\Delta S,}$ (28)
${F=\frac{4\pi Mmc^{3}}{\hbar N}.}$ (29)
On the other side, by substituting Eq. (24) into Eq. (25), we can obtain
${N=\frac{c^{3}}{G\hbar}A\left[1-\frac{\hbar^{2}\beta}{4\pi A}\ln\left(\frac{A}{4l_{p}^{2}}\right)\right].}$ (30)
Substituting Eq. (30) into Eq. (29) gives the modified Newtonian force
${\mathcal{F}=G\frac{Mm}{r^{2}}\left[1+\frac{\hbar^{2}\beta}{16\pi^{2}r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right].}$ (31)
Now consider a spherical system whose mass $M(r)$. The unmodified Newtonian force $dF$ acted on a spherical shell with thickness $dr$ and surface area $dA$ due to the matter inside it writes
${dF=A\>dP=-G\frac{M(r)\>dM(r)}{r^{2}},}$ (32)
where $dM(r)$ denotes the mass of the spherical shell, and $dP$ is the outward pressure balancing the gravitational attraction. Recall $dM(r)=\rho(r)A\>dr$, so
${\frac{1}{\rho(r)}\frac{dP}{dr}=-\frac{GM}{r^{2}}}.$ (33)
=We can easily obtain
${\frac{dP}{dr}=-\frac{GM(r)\rho(r)}{r^{2}}}.$ (34)
On the other side, the Newtonian force within MENG reads
${d\mathcal{F}=A\>dP=-G\frac{M(r)dM(r)}{r^{2}}\left[1+\frac{\hbar^{2}\beta}{16\pi^{2}r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right]}.$ (35)
For this case, Eq. (34) now becomes
${\frac{dP}{dr}=-\frac{GM(r)\rho(r)}{r^{2}}\left[1+\frac{\hbar^{2}\beta}{16\pi^{2}r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right]}.$ (36)
By combining Eq. (36) and Eq. (3), one can obtain the hydrostatic mass of GCs within MENG
${M_{MENG}=-\frac{k_{b}\>T\>r}{G\>\mu\>m_{p}}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)\left[1+\frac{\hbar^{2}\beta}{16\pi^{2}r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right]^{-1}.}$ (37)
@@ 3 Numerical results
Figure 1: The linear fit of the relation between hydrostatic masses and the obseved baryonic mass $M_{bar}$ of 10 GCs. (a) $M_{EiBI}$ with $\kappa=5.80\times 10^{40}$. (b) (c) $M_{BHG}$ with $\Upsilon=1.65\times 10^{71}$.
@@
(kpc) ($\mathscr{M}_{\odot}$) $\beta_{0}=-1.656\times 10^{110}$ $\kappa=5\times 10^{40}$ $\Upsilon=-0.1655\times 10^{69}$
@@
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [19], [43], and [36]. In Ref. [19], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [43], $\kappa=0,2$ could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [19], polytropic star in Ref. [43], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass. However, although $\beta_{0}$ is a free parameter, setting its value to produce a significant impact observable on astronomical scales is not appropriate, as $\beta_{0}$ originates from quantum corrections.
Our result leads to an argument that if the values of the free parameters $\beta_{0}$, $\kappa$, and $\Upsilon$ are closer to zero than the values shown in Table 1, one can safely use the Newtonian calculation for hydrostatic mass of GCs, and the corrections of the gravitational theories used in this paper are negligible. The typical result is also obtained within the formalism which includes GR corrections, see Refs. [3, 16].
@@
The linear fit of the EiBI theory with $\kappa=5.80\times 10^{40}$ gives the slope of $0.126\pm 0.086$, while the BHG with $\Upsilon=1.65\times 10^{71}$ gives the slope of $0.023\pm 0.119$. The linear fit of EiBI theory is shown in Fig. 1(a), while the linear fit of BHG is shown in Fig. 1(b). The slope of the results from EiBI theory is closer to unity compared to that of BHG, indicating that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs, under the same assumption used in Ref. [41] of no dark matter. However, these slope values remain significantly different from unity, suggesting that neither EiBI theory nor BHG fully resolves the mass discrepancy. The exploration of alternative modified gravity theories with appropriate parameters to address the mass discrepancy problem is deferred to future work.
It is important to note that for both EiBI theory and BHG when we set $\kappa$ and $\Upsilon$ as positive numbers, the mass of GCs decreases. As we magnify both $\kappa$ and $\Upsilon$, the hydrostatic mass of GCs increases. Based on this mathematical fact, a trivial thought might arise that we should magnify the values of $\kappa$ and $\Upsilon$ to obtain better linear fits. But in fact, it will not happen, since if we set $\kappa$ and $\Upsilon$ exceed the values used in the linear fits shown by Fig. 1(a) and Fig. 1(b), some masses of GCs reach negative values, which is not physical.